# CAST: Component-Aligned 3D Scene Reconstruction from an RGB Image

### v1

@@ 6.3 Evaluations
Figure 10: Comparison of scene reconstruction with and without relational graph constraints. By integrating relational graph constraints, our method ensures both physical plausibility and accurate alignment with the intended scene, maintaining correct spatial relationships.
@@
To assess the effectiveness of our pose alignment module, we compared it with common pose estimation methods such as Iterative Closest Point (ICP) [2, 6] and differentiable rendering [39]. The generated mesh was provided to different pose estimation methods to align it with the reference RGB image and its corresponding depth prediction. For the ICP method, we uniformly sampled a point cloud from the generated mesh and normalized both the sampled and estimated point clouds by their bounding boxes to address scale differences. We used the ICP implementation in Open3D [95] to register these two normalized point clouds. For differentiable rendering, we optimized the rotation and translation parameters to transform the generated mesh so that the rendered image aligned with the reference RGB image. As shown in Fig. 9, our method surpasses both ICP and differentiable rendering in alignment accuracy. ICP often struggles with accurate pose estimation due to outliers in point clouds, unknown object scales, and symmetrical or repetitive geometries, which can lead to local minima. Differentiable rendering, on the other hand, is significantly impacted by occlusions in the RGB input, disrupting the optimization of object poses and preventing precise alignment with the input image. Our results show that our pose alignment module outperforms traditional ICP and differentiable rendering methods, demonstrating its robustness in accurately estimating object poses from generated meshes and improving alignment with input images.
@@
In CAST, physical constraints are essential for achieving realistic object interactions and maintaining spatial coherence within a scene. While we address common challenges such as occlusions and incomplete views, issues like floating objects, penetration, and misaligned spatial relationships still occur. As shown in Fig. 10, scenes generated without relational constraints may appear physically inconsistent, when only physical simulation is applied, objects adhere to physical laws, but their relative positioning and overall arrangement can differ significantly from the intended scene (e.g., an onion might fall off a surface, disrupting the original composition). By incorporating relational graph constraints, our method ensures that the objects not only comply with physical feasibility but also align with the intended scene layout, preserving both the physical plausibility and the desired spatial relationships.
Table 3: Quantitative ablation study of the MAE module, point cloud conditioning (PCD), and the iterative refinement strategy (iter.). For simplicity, we only display the added key component in each row.
@@
Figure 11: CAST enables realistic physics-based animations, immersive game environments, and efficient real-to-simulation transitions, driving innovation across various fields.
7 Conclusions
@@
The quality of scene generation in CAST is heavily dependent on the underlying object generation model. At present, the model still lacks sufficient detail and precision, this limitation leads to noticeable inconsistencies in the generated objects, affecting their alignment and spatial relationships in the scene. Although additional modules have been incorporated to enhance object robustness and similarity, the need for more advanced and robust generation models remains. A more detailed and accurate object generator could significantly improve the overall scene quality and enhance its real-world applicability.
A notable limitation of the current method is the absence of lighting estimation and background modeling. Without realistic lighting, the interactions between objects and their surroundings may lack natural shading and illumination effects, impacting the visual realism and immersion of the generated 3D environments. Future enhancements in CAST could benefit from integrating advanced techniques for lighting estimation and background modeling, which would significantly enrich the contextual depth and visual fidelity of the scenes.

### v2

@@ 6.3. Evaluations
Figure 10. Comparison of scene reconstruction with and without relational graph constraints. By integrating relational graph constraints, our method ensures both physical plausibility and accurate alignment with the intended scene, maintaining correct spatial relations.
@@
To assess the effectiveness of our pose alignment module, we compared it with common pose estimation methods such as Iterative Closest Point (ICP) (Arun et al., 1987; Best, 1992) and differentiable rendering (Laine et al., 2020). The generated mesh was provided to different pose estimation methods to align it with the reference RGB image and its corresponding depth prediction. For the ICP method, we uniformly sampled a point cloud from the generated mesh and normalized both the sampled and estimated point clouds by their bounding boxes to address scale differences. We used the ICP implementation in Open3D (Zhou et al., 2018) to register these two normalized point clouds. For differentiable rendering, we optimized the rotation and translation parameters to transform the generated mesh so that the rendered image aligned with the reference RGB image. As shown in Fig. 9, our method surpasses both ICP and differentiable rendering in alignment accuracy. ICP often struggles with accurate pose estimation due to outliers in point clouds, unknown object scales, and symmetrical or repetitive geometries, which can lead to local minima. Differentiable rendering, on the other hand, is significantly impacted by occlusions in the RGB input, disrupting the optimization of object poses and preventing precise alignment with the input image. Our results show that our pose alignment module outperforms traditional ICP and differentiable rendering methods, demonstrating its robustness in accurately estimating object poses from generated meshes and improving alignment with input images.
@@
In CAST, physical constraints are essential for achieving realistic object interactions and maintaining spatial coherence within a scene. While we address common challenges such as occlusions and incomplete views, issues like floating objects, penetration, and misaligned spatial relations still occur. As shown in Fig. 10, scenes generated without relational constraints may appear physically inconsistent, when only physical simulation is applied, objects adhere to physical laws, but their relative positioning and overall arrangement can differ significantly from the intended scene (e.g., an onion might fall off a surface, disrupting the original composition). By incorporating relational graph constraints, our method ensures that the objects not only comply with physical feasibility but also align with the intended scene layout, preserving both the physical plausibility and the desired spatial relations.
Table 3. Quantitative ablation study of the MAE module, point cloud conditioning (PCD), and the iterative refinement strategy (iter.). For simplicity, we only display the added key component in each row.
@@
Figure 11. CAST enables realistic physics-based animations, immersive game environments, and efficient real-to-simulation transitions, driving innovation across various fields.
7. Conclusions
@@
Figure 12. In some scenes, transparent glass, textiles, and fabrics are difficult to express, as the mesh struggles to represent them realistically.
@@
The quality of scene generation in CAST is heavily dependent on the underlying object generation model. At present, the model still lacks sufficient detail and precision, this limitation leads to noticeable inconsistencies in the generated objects, affecting their alignment and spatial relations in the scene. Additionally, the current mesh representation struggles with materials like textiles, glasses or fabrics, often appearing unnatural, and fails to accurately depict transparent materials, as shown in Fig. 12. Although additional modules have been incorporated to enhance object robustness and similarity, the need for more advanced and robust generation models remains. A more detailed and accurate object generator could significantly improve the overall scene quality and enhance its real-world applicability.
A notable limitation of the current method is the absence of lighting estimation and background modeling. Without realistic lighting, the interactions between objects and their surroundings may lack natural shading and illumination effects, impacting the visual realism and immersion of the generated 3D environments. For enhanced visual realism, we employ an off-the-shelf panoramic HDR generation tool (Hyper3D, 2025) combined with preset lighting conditions in blender manually. Future enhancements in CAST could benefit from integrating advanced techniques for lighting estimation and background modeling, which would significantly enrich the contextual depth and visual fidelity of the scenes.
@@
Acknowledgements.
This work was supported by National Key R&D Program of China (2022YFF0902301), NSFC programs (61976138, 61977047), STCSM (2015F0203-000-06), and SHMEC (2019-01-07-00-01-E00003). We also acknowledge support from Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI), MoE Key Lab of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), Core Facility Platform of Computer Science and Communication of ShanghaiTech University, and HPC Platform of ShanghaiTech University.
Appendix A GPT-4v prompt
The following prompt instructs GPT-4V to act as an object–relationship analyst for a numbered scene image. It defines six relationship types (Stack, Lean, Hang, Clamped, Contained, Edge/Point), sets strict formatting rules (JSON objects in a list). For full details, see the appendix prompt as below.
Listing 1: System prompt for GPT-4V
prompting_text_system = """
Youareanexpertinobjectrecognitionandspatialreasoning.
###TaskDescription###
Analyzeanimagewithnumberedobjectsanddeterminetheirrelationships.
Foreachpairofrelatedobjects,outputaJSONobjectcontainingtherelationshipdetails.
Ensureyououtputallpossiblerelationships,eventhosethatmaybedifficulttojudgeorlessobvious.
###RelationshipDefinitions###
1.**Stack**:Object1isontopofObject2(Object2supportsObject1frombelow)
2.**Lean**:Object1isleaningagainstObject2(Object2supportsObject1laterally)
3.**Hang**:Object1ishangingfromObject2(Object2supportsObject1fromabove)
4.**Clamped**:Object1isclampedbyObject2(Object2gripsObject1onmultiplesides)
5.**Contained**:Object1isinsideObject2(Object2enclosesObject1)
6.**Edge/Point**:Object1istouchingObject2atanedgeorpoint(minimalcontact,nosignificantsupport)
###ImportantNote###
Onlyobjectsthatareincontactwitheachothershouldhavearelationship.
Foreachrelationship,alwaysensurethefollowing:
1.Usethecorrectrelationshiptype.
2.Provideaclearexplanationoftherelationship.
3.Forcasesthatareincontactbuthardtochoosewhichtype,use"Stack".
###OutputFormat###
␣␣’pair’:[obj1_num,obj2_num],
␣␣’relationship’:’Stack’/’Lean’/’Hang’/’Clamped’/’Contained’/’Edge/Point’,
␣␣’reason’:’explanation’
###Examples###
{’pair’:[1,2],␣␣’relationship’:’Stack’,␣␣␣␣␣␣’reason’:’Book(1)isstackedontopofTable(2)’}
{’pair’:[3,4],␣␣’relationship’:’Lean’,␣␣␣␣␣␣␣’reason’:’Achair(3)isleaningagainstawall(4)’}
{’pair’:[5,6],␣␣’relationship’:’Hang’,␣␣␣␣␣␣␣’reason’:’Alamp(5)ishangingfromtheceiling(6)’}
{’pair’:[7,8],␣␣’relationship’:’Clamped’,␣␣␣␣’reason’:’Apipe(7)isclampedbyabracket(8)’}
{’pair’:[9,10],’relationship’:’Contained’,␣␣’reason’:’Apencil(9)isinsideapencilcase(10)’}
{’pair’:[11,12],’relationship’:’Edge/Point’,’reason’:’Abook(11)andapen(12)aretouchingattheedge’}
Onlyobjectsthatareincontactwitheachothershouldhavearelationship.
content_user = [
"type": "text",
"text": (’’
␣␣␣␣␣␣␣␣␣␣␣␣"###ObjectDetails###"
␣␣␣␣␣␣␣␣␣␣␣␣"IhavelabeledabrightnumericIDatthecenterforeachvisualobjectintheimage."
␣␣␣␣␣␣␣␣␣␣␣␣"PleaseanalyzeallrelationshipsbetweenthenumberedobjectsandoutputJSONobjectsfollowingthespecifiedformat."
␣␣␣␣␣␣␣␣␣␣␣␣"Ensureeachrelationshipincludes:"
␣␣␣␣␣␣␣␣␣␣␣␣"1.Thecorrectrelationshiptype"
␣␣␣␣␣␣␣␣␣␣␣␣"2.Aclearreasonfortherelationship"
␣␣␣␣␣␣␣␣␣␣␣␣"ListallrelationshipsasaJSONarray."
"type": "image_url",
"image_url": {"url": f"data:image/png;base64,{base64_image}"},